# Scaling Laws for Multilingual Language Models

### v1

@@ Abstract
We propose a novel scaling law for general-purpose decoder-only language models (LMs) trained on multilingual data, addressing the problem of balancing languages during multilingual pretraining. A primary challenge in studying multilingual scaling is the difficulty of analyzing individual language performance due to cross-lingual transfer. To address this, we shift the focus from individual languages to language families. We introduce and validate a hypothesis that the test cross-entropy loss for each language family is determined solely by its own sampling ratio, independent of other languages in the mixture. This insight simplifies the complexity of multilingual scaling and make the analysis scalable to an arbitrary number of languages. Building on this hypothesis, we derive a power-law relationship that links performance with dataset size, model size and sampling ratios. This relationship enables us to predict performance across various combinations of the above three quantities, and derive the optimal sampling ratios at different model scales. To demonstrate the effectiveness and accuracy of our proposed scaling law, we perform a large-scale empirical study, training more than 100 models on 23 languages spanning 5 language families. Our experiments show that the optimal sampling ratios derived from small models (85M parameters) generalize effectively to models that are several orders of magnitude larger (1.2B parameters), offering a resource-efficient approach for multilingual LM training at scale.
@@ 1 Introduction
Although there exist some studies that examine multilingual scaling laws (Fernandes et al., 2023; Chen et al., 2024), they often focus on the specific problem of neural machine translation (NMT). These studies predominantly utilize encoder-decoder transformer architectures, which are not widely applied to generation tasks. Furthermore, they typically restrict their analysis to bilingual settings without capturing the complexities of language interactions. This omission is critical, as cross-lingual transfer, a key benefit in multilingual training, plays a significant role in improving performance across languages (Arivazhagan et al., 2019; Xian et al., 2022; Patra et al., 2023).
@@
Figure 1: We propose a multilingual scaling law connecting the test cross-entropy loss ($\mathcal{L}$) with model size ($N$), dataset size ($D$) and sampling ratios for different language families (${\bm{p}}$). The plots illustrate a power-law relationship by varying one quantity while fixing the other two for five language families.
@@ 3.2 Hypothesis
Hypothesis testing. We use a controlled experiment to test our hypothesis. We train an LM on a data mixture containing three language families: Romance, Indic and Slavic. In the training, we fix the sampling ratio of Romance to be either 0.2 or 0.5, and vary the ratios of the other two families. In Figure 2 (left), we see that with a fixed sampling ratio, the Romance loss does not noticeably change regardless of the combination ratios of Indic and Slavic data. The stability of the Romance loss suggests that the primary source of performance variation is the sampling ratio of the family itself, rather than interference or assistance from other language families. This observation aligns with our hypothesis that cross-family transfer is minimal, making it feasible to model each language family’s loss as a function of its own sampling ratio. We also perform the same experiment with other family combinations, detailed in Appendix C.
Figure 2: Left: The Romance loss remains stable as the Indic sampling ratio varies, indicating minimal cross-family transfer and supporting our hypothesis that each language family’s performance is primarily influenced by its own sampling ratio. Right: In contrast, when both groups contain Indic languages, Group 1 loss decreases as Group 2 sampling ratio increases, demonstrating significant cross-group transfer. This underscores the importance of grouping by language families for accurate analysis. The loss values are normalized by the mean loss at $p=0.2$ to align the plot scales. Shaded areas indicate standard deviation.
@@ 4.1 Experimental setup
$p_{i}=\frac{q_{i}^{\alpha}}{\sum_{i=1}^{n}q_{j}^{\alpha}}\text{ with }q_{i}=\frac{D_{i}}{D}.$
@@ Appendix C Additional experiments
Figure 6: The Sino loss does not vary with proportions of other jointly trained family.
We repeat the hypothesis validation experiments as in Section 3.2 with another language family combination: Sino-Tibetan, Slavic and Germanic. Here, we fix the sampling ratio of Sino-Tibetan to be either 0.2 or 0.5, and vary the ratios of the other two families. From Figure 6, we still see that the hypothesis holds well, as the Sino-Tibetan loss does not noticeably change when varying sampling ratios of other families. The result further validates 1.

### v2

@@ Abstract
We propose a novel scaling law for general-purpose decoder-only language models (LMs) trained on multilingual data, tackling the problem of balancing languages during multilingual pretraining. A primary challenge in studying multilingual scaling is the difficulty of analyzing individual language performance due to cross-lingual transfer. To address this, we shift the focus from individual languages to language families. We introduce and validate a hypothesis that the test cross-entropy loss for each language family is determined solely by its own sampling ratio, independent of other languages in the mixture. This insight simplifies the complexity of multilingual scaling and make the analysis scalable to an arbitrary number of languages. Building on this hypothesis, we derive a power-law relationship that links performance with dataset size, model size and sampling ratios. This relationship enables us to predict performance across various combinations of the above three quantities, and derive the optimal sampling ratios at different model scales. To demonstrate the effectiveness and accuracy of our proposed scaling law, we perform a large-scale empirical study, training more than 100 models on 23 languages spanning 5 language families. Our experiments show that the optimal sampling ratios derived from small models (85M parameters) generalize effectively to models that are several orders of magnitude larger (1.2B parameters), offering a resource-efficient approach for multilingual LM training at scale.
@@ 1 Introduction
Although there exist some studies that examine multilingual scaling laws (Gordon et al., 2021; Ghorbani et al., 2022; Fernandes et al., 2023; Chen et al., 2024), they often focus on the specific problem of neural machine translation (NMT). These studies predominantly utilize encoder-decoder transformer architectures, which are not widely applied to generation tasks. Furthermore, they typically restrict their analysis to bilingual settings without capturing the complexities of language interactions. This omission is critical, as cross-lingual transfer, a key benefit in multilingual training, plays a significant role in improving performance across languages (Arivazhagan et al., 2019; Xian et al., 2022; Patra et al., 2023).
@@
Figure 1: We propose a multilingual scaling law connecting the test cross-entropy loss ($\mathcal{L}$) with model size in number of parameters ($N$), dataset size ($D$) and sampling ratios for different language families (${\bm{p}}$). The plots illustrate a power-law relationship by varying one quantity while fixing the other two for five language families.
@@ 3.2 Hypothesis
Hypothesis testing. We use a controlled experiment to test our hypothesis. We train an LM on a data mixture containing three language families: Romance, Germanic and Slavic. In the training, we fix the sampling ratio of Romance to be either 0.2 or 0.5, and vary the ratios of the other two families. At the individual language level, one might expect high-resource languages like English (within the Germanic family) to transfer knowledge to related languages in the Romance family (e.g., Spanish, French). However, our hypothesis focuses on cross-family transfer instead of cross-lingual transfer among individual languages. In Figure 2 (left), we observe that with a fixed sampling ratio, the Romance loss does not noticeably change regardless of the combination ratios of Germanic and Slavic data. This stability indicates that the primary source of performance variation is the sampling ratio of the family itself, rather than cross-family interactions. This finding supports our hypothesis that cross-family transfer is minimal, making it feasible to model each language family’s loss as a function of its own sampling ratio. To strengthen our findings, we conduct the same experiment with three additional family combinations, detailed in Appendix C.
Figure 2: Left: The Romance loss remains stable as the Germanic sampling ratio varies, indicating minimal cross-family transfer and supporting our hypothesis that each language family’s performance is primarily influenced by its own sampling ratio. Right: In contrast, when both groups contain Indic languages, Group 1 loss decreases as Group 2 sampling ratio increases, demonstrating significant cross-group transfer. This underscores the importance of grouping by language families for accurate analysis. The loss values are normalized by the mean loss at $p=0.2$ to align the plot scales. Shaded areas indicate standard deviation.
@@ 4.1 Experimental setup
$p_{i}=\frac{q_{i}^{\alpha}}{\sum_{j=1}^{n}q_{j}^{\alpha}}\text{ with }q_{i}=\frac{D_{i}}{D}.$
@@ Appendix C Additional experiments
Figure 6: Further validation on three family combinations: {Romance, Indic, Slavic} (left), {Germanic, Romance, Slavic} (middle), {Sino-Tibetan, Germanic, Slavic} (right). In all three cases, the loss for the family with a fixed ratio does not vary with proportions of other jointly trained family.
We repeat the hypothesis validation experiments as in Section 3.2 with three other language family configurations: {Romance, Indic, Slavic}, {Germanic, Romance, Slavic}, {Sino-Tibetan, Germanic, Slavic}, where the underlined language family has a fixed sampling ratio of either 0.2 or 0.5, and the other two families have varying ratios. From Figure 6, we still see that in all three cases, the hypothesis holds well, as the loss of the fixed family does not noticeably change when varying sampling ratios of other families. The result further validates 1.